# Retrieval-Generation Alignment for End-to-End Task-Oriented Dialogue System

### v1

@@ 3.4.1 Prefix
In this approach, we create a mapping function that assigns special tokens representing meta knowledge to each entity. For instance, an entity ranked second in retrieval order, with middle retrieval confidence, and not yet mentioned in the context would be mapped to the set of <2nd-entity>, <mid-confidence>, <new-entity>222The complete mapping is provided in Appendix B.1. These prefix tokens are then concatenated with the corresponding entity and input to the generator during both training and inference stages.
@@ 3.4.2 Prompt
To fully leverage the generator’s language modeling capability, we explore using prompts to incorporate meta knowledge. Here, we design a mapping function that assigns each entity a set of prompts, which are natural language sentences representing the meta knowledge. For example, a prompt can be “This is the top-1 recalled entity with low confidence”.333The complete mapping is provided in Appendix B.2 Similar to the prefix approach, these prompts are concatenated with the corresponding entities and fed into the generator.
@@ 4.1 Dataset
We evaluate our MK-TOD on three task-oriented dialogue datasets: MultiWOZ 2.1 (MWOZ) (Eric et al., 2020), CamRest (Wen et al., 2017), and Stanford Multi-Domain (SMD) (Eric et al., 2017). We compare the methods with two different settings about the knowledge bases: First, each dialogue has a small session-level knowledge base associated with the user goal, which is referred to as the condensed knowledge base. Second, Q-TOD (Tian et al., 2022) proposes to construct a dataset-level large-scale knowledge base for MWOZ and CamRest by injecting all the condensed knowledge bases. There are 223 and 112 entities in the large-scale knowledge base for MWOZ and CamRest, respectively. The setting of the large-scale knowledge base imposes more challenges for the retrieval and utilization of knowledge. Other detailed statistics of these datasets are shown in Appendix A.
@@
4.2 Implementation Details
Consistent with previous studies (Qin et al., 2019), we initialize the retriever through pre-training with distant supervision to prevent collapsed representations. Additional details on the pre-training process can be found in Appendix D.
@@ 4.2 Implementation Details
| Ours (Large) | 16.69 | 53.59 | 84.58 | 27.32 | 72.77 | 87.05 |
| Ours (Large) | 17.15 | 52.99 | 88.66 | 26.88 | 72.92 | 92.94 |
| Ours (Large) | 17.40 | 53.26 | 93.19 | 27.82 | 71.98 | 92.70 |
@@ 4.3 Baseline Methods
Implicit retrieval: These methods combine knowledge retrieval and response generation in a single model, including KB-Retriever (Qin et al., 2019), GLMP (Wu et al., 2019), DF-Net (Qin et al., 2020), EER (He et al., 2020b), FG2Seq (He et al., 2020a), CDNET (Raghu et al., 2021), GPT-KE (Madotto et al., 2020) and ECO (Huang et al., 2022).
@@ 5 Results and Analysis
5.1 Overall Results with Large-Scale KBs
We observe that Q-TOD’s retriever outperforms ours due to their utilization of an additional query generator, which helps generate more accurate retrieval queries. However, even when employing T5-Base and a relatively weaker retriever, our method still surpasses Q-TOD in terms of BLEU and Entity F1 by a significant margin. This indicates that our proposed method effectively utilizes the retrieved knowledge better than Q-TOD.
@@ 5.4.2 Types of Meta Knowledge
We compare different types of meta knowledge, and the results are presented in Table 4. The findings indicate that using a single type of meta knowledge yields inferior performance compared to combining all three types. Furthermore, an interesting observation emerges when using the prefix: the retrieval order outperforms other types of meta knowledge. In contrast, when using the prompt, the results are reversed. We attribute this phenomenon to the design of the prefix and prompt. Representing meta knowledge with a prefix introduces a higher diversity in ranking order since a distinct prefix is assigned to each ranking order. This increased diversity enables the generator to better distinguish the recalled entities. On the other hand, the distinction between retrieval confidence and co-occurrence in the prefix setting is less obvious. In contrast, when representing meta knowledge with a prompt, the retrieval order becomes less diverse.
@@ 5.5 Behavior of Generator
We examine how the generator utilizes the retrieved knowledge with the assistance of meta knowledge on the MWOZ test set. For our model and the baseline, which is T5-Large, we gather all their responses that contain entities, and analyze the percentage of retrieved entities that correctly appear in the responses according to retrieval order and confidence. As illustrated in Figure 4, our generator exhibits a higher propensity than the baseline to utilize entities with both a high retrieval order and high confidence. We have assessed the retrieval results on the MWOZ test set. The findings demonstrate that our retriever adeptly recalls 80.69% of the gold entities with a top-1 retrieval order, which directly correlates with the system-wide performance enhancement. This observation suggests that our proposed meta knowledge aids the generator in developing an inductive bias to prioritize entities that are highlighted by the retriever.

### v2

@@ 3.4.1 Prefix
In this approach, we create a mapping function that assigns special tokens representing meta knowledge to each entity. For instance, an entity ranked second in retrieval order, with middle retrieval confidence, and not yet mentioned in the context would be mapped to the set of <2nd-entity>, <mid-confidence>, <new-entity>222The complete mapping is discussed in Appendix B.1. These prefix tokens are then concatenated with the corresponding entity and input to the generator during both training and inference stages.
@@ 3.4.2 Prompt
To fully leverage the generator’s language modeling capability, we explore using prompts to incorporate meta knowledge. Here, we design a mapping function that assigns each entity a set of prompts, which are natural language sentences representing the meta knowledge. For example, a prompt can be “This is the top-1 recalled entity with low confidence”.333The complete mapping is discussed in Appendix B.2 Similar to the prefix approach, these prompts are concatenated with the corresponding entities and fed into the generator.
@@ 3.6 Model Inference
3.7 Discussion
The conceptual of meta-knowledge is to empower the generator with the generator’s capability to leverage retrieved information to enhance the generation process. Noteworthy is the fact that, while we deliberately furnish the generator with meta-knowledge such as retrieval order and confidence, this does not obligate the generator to strictly use the higher-ranking entities. Instead, our approach encourages the generators to autonomously differentiate between these entities. Specifically, for T5, we integrate meta-knowledge as a part of the model input, facilitating its assimilation during the training process. With respect to ChatGPT, our method capitalizes on the model’s pre-existing knowledge to comprehend the essence of meta-knowledge. The results presented in Sections 5.5 exemplify the behaviors acquired by the generator and how these behaviors contribute to enhanced performance.
@@ 4.1 Dataset
We evaluate our MK-TOD on three task-oriented dialogue datasets: MultiWOZ 2.1 (MWOZ) (Eric et al., 2020), CamRest (Wen et al., 2017), and Stanford Multi-Domain (SMD) (Eric et al., 2017). We compare the methods with two different settings about the knowledge bases: First, each dialogue has a small session-level knowledge base associated with the user goal, which is the typical setting of previous work. Second, all conversations shares a dataset-level large-scale knowledge base by injecting all the session-level knowledge bases. There are 223 and 112 entities in the large-scale knowledge base for MWOZ and CamRest, respectively. Other detailed statistics of these datasets are shown in Appendix A.
@@
| Ours (Large) | 16.69 | 53.59 | 92.93 | 27.32 | 72.77 | 91.08 |
| Ours (Large) | 17.15 | 52.99 | 94.42 | 26.88 | 72.92 | 95.41 |
| Ours (Large) | 17.40 | 53.26 | 95.22 | 27.82 | 71.98 | 95.38 |
@@ 4.3 Baseline Methods
Implicit retrieval: These methods combine knowledge retrieval and response generation in a single model, including DF-Net (Qin et al., 2020), EER (He et al., 2020b), FG2Seq (He et al., 2020a), CDNET (Raghu et al., 2021) and GPT-KE (Madotto et al., 2020).
@@ 5 Results and Analysis
5.1 Overall Results with Large-Scale KBs
Overall, MK-TOD outperforms all previous methods with the same scale of generator model, indicating the effect of our proposed meta-knowledge. Further more, Q-TOD’s retriever can achieve comparable and even higher performance than ours due to their utilization of an additional query generator. However, even when employing T5-Base and a relatively weaker retriever, our method still surpasses Q-TOD in terms of BLEU and Entity F1 by a significant margin. This indicates that our proposed method effectively utilizes the retrieved knowledge better than Q-TOD. It is also noteworthy that the Entity F1 score of T5-Large in CamRest is poorer than that of T5-Base, we attribute this to the small size of CamRest’s training data, which includes merely 406 dialogues, leading to overfitting particularly for larger models.
@@ 5.4.2 Types of Meta Knowledge
We compare different types of meta knowledge, and the results are presented in Table 4. The findings indicate that using a single type of meta knowledge yields inferior performance compared to combining all three types. Furthermore, an interesting observation emerges when using the prefix: the retrieval order outperforms other types of meta knowledge. In contrast, when using the prompt, the results are reversed. We attribute this phenomenon to the design of the prefix and prompt. Representing meta knowledge with a prefix introduces a higher diversity in ranking order since a distinct prefix is assigned to each ranking order. This increased diversity enables the generator to better distinguish the recalled entities. On the other hand, the distinction between retrieval confidence and co-occurrence in the prefix setting is less obvious. In contrast, when representing meta knowledge with a prompt, the retrieval order becomes less diverse, since only the numbers representing the retrieval order are varied.
@@ 5.5 Behavior of Generator
We examine how the generator utilizes the retrieved knowledge with the assistance of meta knowledge on the MWOZ test set. For our model and the baseline, which is T5-Large, we gather all their responses that contain entities, and analyze the percentage of retrieved entities that appear in the responses according to retrieval order and confidence. As illustrated in Figure 4, our generator exhibits a higher propensity than the baseline to utilize entities with both a high retrieval order and high confidence. Further more, We have assessed the retrieval results on the MWOZ test set. The findings demonstrate that our retriever adeptly recalls 80.69% of the gold entities with a top-1 retrieval order, which directly correlates with the system-wide performance enhancement. This observation suggests that our proposed meta knowledge aids the generator in developing an inductive bias to prioritize entities that are highlighted by the retriever.
@@ B.2 Prompt of Meta Knowledge
B.3 Discussion
How generator learn to distinguish entities with meta knowledge?
Our method aims to enable the models to comprehend meta knowledge, such as retrieval order and confidence scores, and then compare the retrieved entities. To achieve this, we employ distinct strategies based on whether the models are fine-tunable or not. For fine-tunable models like T5, we include the meta knowledge as a part of the model input, enabling its assimilation during the training process. In the case of non-fine-tunable models like ChatGPT, our approach utilizes the model’s existing knowledge to grasp the essence of meta knowledge. This is implemented by providing an explanation of meta knowledge within the system prompt, as elaborated in Appendix B.1 and Appendix B.2.
How are the prompts of meta knowledge designed?
When it comes to prompts for T5 models, given their fine-tuning capability, our emphasis lies in the effectiveness of implementation with the pytorch code, and we leave the comprehension of prompts to the model itself during the training phrase. To achieve this, we craft the prompts in a manner that permits them to be tokenized into sequences of uniform length with the tokenizer, which facilitate the efficient mapping to entities through the torch.gather operation.
For prompts for ChatGPT, the final prompts are yield thought several trials and evaluations on the valid set of MWOZ.